nowledge Check 01 On January 1, Year 1, the Starshina Company paid $25,000 for a photocopier. The company estimates that the photocopies will produce 1,000,000 copies and will have an estimated residual value of $5.000 The company used to photocopier to make 220.000 copies during Year 1. The company uses the units-of-production method. What is the amount of depreciation expense for Year 1?
Knowledge Check 01 On January 1. Year 1, the Starshina Company paid $25,000 for a photocopier with an estimated useful life of 4 years, and an estimated residual value of $5,000. The company uses the double-declining-balance method What is the amount of depreciation expense for Year 3? eciation expense
Answer #1
Ans.1 Straight line depreciation   = (Cost - Residual value) / Life in years
(25000 - 5000) / 4
5000
In stratight line method the depreciation for each year will be equal.
Ans.2 Units of production method = ( Cost - residual value) / Estimated production * Actual production
(25000 - 5000) / 1000000 * 220000
20000 / 1000000 * 220000
4400
Ans.3 Double declining method:
Double declining balance rate   =   2 * 1 / life of asset
2 * 1 / 4
0.5
Depreciation = Remaining value at the beg. Of year * 0.50
Year Remaining value at beg. Of year (a) Depreciation (b = a*0.50) Remaining value at end Of year (a-b)
1 25000 12500 12500
2 12500 6250 6250
3 6250 1250 5000
The remaining value at the end of year can not fall below the residual value
the actual depreciation for year 3 is 6250 * 0.50 = 3125 that would result the
remaining value at the year end = 6250 - 3125 = 3125 which is less than the
residual value, which is not possible
so we will calculate the accurate depreciation by subtracting the residual value from the beginning value. (6250 - 5000 = 1250)